OpenStudy (anonymous):

The width and length of a rectangle (in feet)are consecutive odd integers. If the length is increased by 5 feet, the area of the resulting rectangle is 60 square feet. What is the area of the original rectangle?

OpenStudy (solomonzelman):

60 = 30 × 2 = 15 × 4 = 12 × 5 = 10 × 6. 12 is the width, so that when it is decreased by 5 you get 2 consecutive odd integers.

OpenStudy (solomonzelman):

And so `7 × 5` is the area of the original shape. Understand ?

OpenStudy (anonymous):

Yes! thank you

OpenStudy (solomonzelman):

Anytime ... this is a help site ! Have fun !